Greenhab Report – April 20th

Crew 211 GreenHab Report – 20-APR-2019

GreenHab Officer: Chloé Peduzzi

Environmental control: Ambient with the door open

Shade cloth: 80%

Average temperatures:

Low: 14°C

High: 43°C

Hours of supplemental light: 0

Daily water usage for crops: 2 gallons

Water in Blue Tank: 125 gallons

Time(s) of watering for crops: 19:30

Changes to crops: Crew 211 has just arrived today

Narrative: Crew 211 has just arrived today

Harvest: None

Support/supplies needed: None

Operations Report – April 20th

Crew 211 Operations Report 20-APR-2019

SOL: /

Name of person filing report: Julien Amalaberque

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run last night): 12
On (last night): 7 pm
Off (this morning): 7 am

Solar— SOC 79 % at 7 pm

Diesel Reading – 25 %

Propane Reading – 49 %

Ethanol Free Gasoline – 4 gallons.

Water (auxiliary tank) – 0 gallons

Water (static tank) – 475 gallons

Auxiliary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – 125 gallons

Water (loft) – Static to Loft Pump used – yes

Water Meter: 9 units

Toilet tank emptied: no

Deimos rover used: No
Hours:
Beginning charge:
Ending charge:
Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, yes
Hours: 51.1h
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Spirit rover used: no
Hours: 87.6h
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Opportunity rover used: no
Hours: 56.8h
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Curiosity rover used: no
Hours: 89.1h
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) none

Reason for use:

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? No

CrewCar used and why, where? Yes, the trip from Grand Junction to MDRS

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: No operation, Chloe will send the first GreenHab report today

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: None, thank you.

Sol Summary – April 20th

Crew 211 Sol Summary Report 20-APR-2019

Sol: /

Summary Title: Discovering the MDRS

Author’s name: Eléonore Lieffrig

Mission Status: Nominal

Sol Activity Summary:

Today was a day of traveling and discovering the MDRS. Five members of the crew arrived at the station at about 1:30 pm. As Chloé, Julien and I discovered our new home, Carl-Henrik and Simon went back to Grand Junction in order to take the rest of the crew and give back our rental car. At the moment (7 pm), they are still on their way.

We took our marks here, cleaned a bit the Hab, chose our bedrooms and got to know the place. Shannon gave us a brief tour of the station, but a longer training is planned with the whole crew tomorrow morning at 10 am.

Tonight, we will read the Hab manual, welcome the rest of the team, and cook ourselves our first tasty Martian dinner.

For the crew 211,

Eléonore Lieffrig, astronomer

Anomalies in work: No problems.

Weather: Warm and windy

Crew Physical Status: Good.

EVA: No EVA today.

Reports to be filed: GreenHab Report, Operations report, Sol Summary Report

Support Requested: None